OpenStudy (anonymous):

how do you simplify the square root of 56x to the fifth

OpenStudy (anonymous):

It comes out to something along the very neat lines of: \[ 6272\sqrt{14x^5} \]Which... is not neat at all.

OpenStudy (anonymous):

how did you do it though????

OpenStudy (anonymous):

Here: \[ \sqrt{56x}=\sqrt{4\cdot14x}=2\sqrt{14x} \]We raise it to the fifth power, to find: \[ \left(2\sqrt{14x}\right)^5=2^5\left((14x)^\frac{1}{2}\right)^5=2^5(14x)^\frac{5}{2}=\\ 2^5(14x)^2(14x)^\frac{1}{2}=2^514^2x^2\sqrt{14x}=6272x^2\sqrt{14x}=\\ 6272\sqrt{14x^5} \]That's the most elaborate approach, you should just try to find stuff, it's fairly easy to get, just requires practice.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i still need help though to whoever can!!